OpenStudy (erinkb99):

EASY CHECK MY ANSWER! The table below gives selected values for the function f(x). Use a trapezoidal estimation, with 6 trapezoids to approximate the value of the integral of f(x) from 1 to 2.

OpenStudy (erinkb99):

USE THIS TABLE x 1 1.1 1.3 1.6 1.7 1.8 2.0 f(x) 1 3 5 8 10 11 14

OpenStudy (erinkb99):

My answer: delta x= (b-a)/n delta x= (2-1)/6 = 1/6 (1/6)/2 (f(1) + 2f(1.1) + 2f(1.3)+2f(1.6)+2f(1.7)+2f(1.8)+f(2.0)) 1/12 (1 + 6 +10 +16+ 20+22+14) 1/12 (89) 89/12 =7.417

OpenStudy (erinkb99):

@Hero @pooja195

OpenStudy (agent0smith):

That looks about right, but i prefer to just draw these and do them geometrically (work out the areas of the trapezoids).

OpenStudy (agent0smith):

btw your answer will be slightly off, because the "height" of each trap. is not 1/6 (the heights vary). Another good reason to just draw it out and do it that way.

OpenStudy (mww):

That formula for Trapezoidal Rule works only works if the width of your subintervals (i.e. space between two adjacent x-values) are regular.

OpenStudy (agent0smith):

@mww that's what my last post is saying